An engineer designs at least one robot a day for 30 days. If a total of 45

robots have been designed, then show that there must have been a series of consecutive

days when exactly 14 robots were designed.


1
Expert's answer
2021-12-31T02:25:06-0500

Let "a_i" be the number of robots designed on the "i"th day. Then we have 30 numbers "a_1,a_2,\u2026,a_{30}" and "a_i\geq 1" for all "1\leq i\leq 30" .

Let "s_k=\sum_{i=1}^ka_i" , it is the number of robots designed on or before the "k"th day.

We have a sequence of 30 numbers "s_1,s_2,…,s_{30},\ \text{where }" "1\leq s_1<s_2<…<s_{30}=45" .

Let us consider new sequence: "s_1+14, \ s_2+14,\ …,\ s_{30}+14" . There are 30 numbers and "s_1+14<s_2+14<\u2026<s_{30}+14=45+14=59" .

There are totally 60 numbers: "s_1,s_2,\u2026,s_{30},s_1+14,s_2+14,\u2026,s_{30}+14" and all of them are less or equal than "59" ("s_i\leq 45" and "s_i+14\leq 59" ).

By the Pigeonhole Principle, at least two of these numbers are equal.

Since "s_i\neq s_j" and "s_i+14\neq s_j+14" for all "i\neq j", it follows that "s_i=s_j+14" for some "i" and "j" .

Now we can conclude that exactly 14 robots were designed from day "j+1" to day "i".